How many atoms of beryllium are in 1.00 g?
1
Expert's answer
2020-09-14T08:25:25-0400

Atomic weight of Beryllium "=9.012g"

The molar mass of Beryllium is therefore"9.012g\/mol"

Converting mass"\to" moles

"1gBe\times" "1molBe\over9.012gBe" "=0.1109molBe"

Moles"\to" Atoms(Multiplying with the Avogadro Constant)

"0.1109molBe\times" "6.022\times10^{23} atomsBe\over1mol Be"

"=6.678\times 10^{22} atoms Be"